### v1

# Calculation of harmonic oscillator brackets in SU(3) basis

Abstract
We present a new approach for the Talmi-Moshinsky transformation representation in the harmonic oscillator basis. We utilize the SU(3) scheme for calculation of harmonic oscillator brackets. Using this scheme we obtain the explicit relations for numeric evaluation and present a computational approach.
keywords:
mathematical methods in physics , algebraic methods , nuclear shell model
PACS:
03.65.Fd 21.60.Cs
1 Introduction
The harmonic oscillator basis (HO) is widely used for the approximation of the wavefunction of the nuclear systems [1, 2]. Traditional ab initio approach for constructing the model Hamiltonian depend on one-particle coordinates. As a result, the model wavefunctions are not translationally invariant by default. Hence, the wavefunction of such a system cannot be represented properly as the center of mass (CM) coordinate of the nucleus has to be eliminated. An additional procedure is required to solve this problem. However, this issue can be avoided by the direct construction of the antisymmetric translationally invariant wavefunction. In this approach, we use the intrinsic coordinates, particularly, the normalized Jacobi coordinates, by transitioning from the one-particle to Jacobi coordinates you get the explicit removal of the CM coordinate [3, 4].
While antisymmetrizing the wavefunction one has to act upon the Jacobi coordinates thus it is beneficial that the different sets of coordinates are related to each other by the orthogonal transformations. The HO basis allows to represent these transformations with a finite number of terms. The key aspect of this approach is the Talmi-Moshinsky transformation and the corresponding brackets (HOB) in the HO basis [5].
HOBs are frequently used in nuclear calculations. So, it is advantageous to have an efficient computational approach for the HOB calculation and there is a notable effort to produce new aproaches for this problem [6, 7, 8].
In our previous papers, we devoted a significant effort to research HOBs. Particularly, we presented the expressions of HO transformations for two and three Jacobi coordinates, analyzed their symmetries, simplifications for special cases of mass-ratio parameters [9, 10]. Also, a set of computer codes were provided [11, 12, 13, 14, 15].
In this paper, we will present a different approach for HOB calculation.
2 Definitions and motivation: HOB in SO(3) basis
In the nuclear structure calculations, the single particle wavefunctions are usually expanded in the HO eigenfunctions. Let us define the HO function as
$\left\langle\boldsymbol{r}|elm\right\rangle=\phi_{elm}\left(\boldsymbol{r}\right)=R_{el}(|r|)Y_{em}(\theta\phi),$ (1)
where $R_{el}(|r|)$ is the radial part and $Y_{em}(\theta\phi)$ is the spherical harmonics. One of the possible function of HO is
$\phi_{elm}(\boldsymbol{r})=\left(-1\right)^{n}\left(\frac{2(n!)}{\Gamma(n+l+\frac{3}{2})}\right)^{1/2}\exp\left(-\frac{r^{2}}{2}\right)r^{l}L_{n}^{(l+\frac{1}{2})}\left(r^{2}\right)Y_{lm}(\frac{\boldsymbol{r}}{r}).$ (2)
In the eq.(2) the corresponding dimensionless eigenvalue equals $(e+\frac{3}{2})$; the principal quantum number $n=\frac{(e-l)}{2}=0,1,2,...$, where $e=2n+l$ and $l$ is the angular momentum. $\Gamma$ denotes the gamma function, $L_{n}^{(l+\frac{1}{2})}$is generalized Laguerre polynomial, and $Y$ is the spherical harmonics. We define the two-particle wavefunction with bound momentum as
$\{\phi_{e_{1}l_{1}}(\boldsymbol{r}_{1})\otimes\phi_{e_{2}l_{2}}(\boldsymbol{r}_{2})\}_{LM}=\left\langle\boldsymbol{r}_{1},\boldsymbol{r}_{2}|e_{1}l_{1},e_{2}l_{2}:LM\right\rangle$
$={\textstyle\sum\limits_{m_{1}m_{2}}}\langle l_{1}m_{1},l_{2}m_{2}|LM\rangle\phi_{e_{1}l_{1}m_{1}}(\boldsymbol{r}_{1})\phi_{e_{2}l_{2}m_{2}}(\boldsymbol{r}_{2}),$ (3)
where $\langle l_{1}m_{1},l_{2}m_{2}|LM\rangle$ is the $SU(2)$ Clebsch-Gordan coefficient. The transformation bracket for two HOs is defined as
$\left|e_{1}l_{1},e_{2}l_{2}:LM\right\rangle={\displaystyle\sum\limits_{e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}}}\left\langle e_{1}l_{1},e_{2}l_{2}:L|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:L\right\rangle_{d}\left|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:LM\right\rangle$ (4)
in [5]. Here $d$ is a nonnegative real number. The variables in the brackets are arranged in a fixed way and the orthogonal transformation of the coordinates $\boldsymbol{r}_{1},\boldsymbol{r}_{2}$ is written as the two-dimensional matrix
$\langle\boldsymbol{r}_{1}^{\prime},\boldsymbol{r}_{2}^{\prime}|\ \boldsymbol{r}_{1},\boldsymbol{r}_{2}\rangle_{d}=\left(\begin{tabular}[]{ll}$\sqrt{\frac{d}{1+d}}$&$\phantom{-}\sqrt{\frac{1}{1+d}}$\\ $\sqrt{\frac{1}{1+d}}$&$-\sqrt{\frac{d}{1+d}}$\end{tabular}\right).$ (5)
This two-dimensional matrix is called the Talmi-Moshinsky transformation. It is parametrized by the variable $d.$ In [5] the original expression for HOBs derived by B.Buck and A.C.Merchant [16]
$\left\langle EL,el:\Lambda\right.\left|e_{1}l_{1},e_{2}l_{2}:\Lambda\right\rangle_{d}$
$=i^{-\left(l_{1}+l_{2}+L+l\right)}\times 2^{-\left(l_{1}+l_{2}+L+l\right)/4}$
×$\times\sqrt{\left(n_{1}\right)!\left(n_{2}\right)!\left(N\right)!\left(n\right)!\left[2\left(n_{1}+l_{1}\right)+1\right]!!\left[2\left(n_{2}+l_{2}\right)+1\right]!!\left[2\left(N+L\right)+1\right]!!\left[2\left(n+l\right)+1\right]!!}$
×$\times\sum_{abcdl_{a}l_{b}l_{c}l_{d}}\left(-1\right)^{l_{a}+l_{b}+l_{c}}2^{\left(l_{a}+l_{b}+l_{c}+l_{d}\right)/2}d^{\left(2a+l_{a}+2d+l_{d}\right)/2}\left(1+d\right)^{-\left(2a+l_{a}+2b+l_{b}+2c+l_{c}+2d+l_{d}\right)/2}$
×$\times\frac{\left[\left(2l_{a}+1\right)\left(2l_{b}+1\right)\left(2l_{c}+1\right)\left(2l_{d}+1\right)\right]}{a!b!c!d!\left[2\left(a+l_{a}\right)+1\right]!!\left[2\left(b+l_{b}\right)+1\right]!!\left[2\left(c+l_{c}\right)+1\right]!!\left[2\left(d+l_{d}\right)+1\right]!!}$
×$\times\left\{\begin{array}[c]{ccc}l_{a}&l_{b}&l_{1}\\ l_{c}&l_{d}&l_{2}\\ L&l&\Lambda\end{array}\right\}\left\langle l_{a}0l_{c}0\right|\left.L0\right\rangle\left\langle l_{b}0l_{d}0\right|\left.l0\right\rangle\left\langle l_{a}0l_{b}0\right|\left.l_{1}0\right\rangle\left\langle l_{c}0l_{d}0\right|\left.l_{2}0\right\rangle$ (9)
was rewritten to make it more computationally feasible and symmetric. The result took the shape of
$\left\langle EL,el:\Lambda\right.\left|e_{1}l_{1},e_{2}l_{2}:\Lambda\right\rangle_{d}$
$=d^{\left(e_{1}-e\right)/2}\left(1+d\right)^{-\left(e_{1}+e_{2}\right)/2}\sum_{e_{a}l_{a}e_{b}l_{b}e_{c}l_{c}e_{d}l_{d}}\left(-d\right)^{e_{d}}\left\{\begin{array}[c]{ccc}l_{a}&l_{b}&l_{1}\\ l_{c}&l_{d}&l_{2}\\ L&l&\Lambda\end{array}\right\}$ (13)
×$\times G\left(e_{1}l_{1};e_{a}l_{a},e_{b}l_{b}\right)G\left(e_{2}l_{2};e_{c}l_{c},e_{d}l_{d}\right)G\left(EL;e_{a}l_{a},e_{c}l_{c}\right)G\left(el;e_{b}l_{b},e_{d}l_{d}\right).$ (14)
$G\left(e_{1}l_{1};e_{a}l_{a},e_{b}l_{b}\right)$ is expressed as trinomial coefficients
$G\left(e_{1}l_{1};e_{a}l_{a},e_{b}l_{b}\right)$ $=\sqrt{\left(2l_{a}+1\right)\left(2l_{b}+1\right)}\left\langle l_{a}0l_{b}0\right|\left.l_{1}0\right\rangle$
×$\times\left[\left(\begin{array}[c]{c}e_{1}-l_{1}\\ e_{a}-l_{a};e_{b}-l_{b}\end{array}\right)\left(\begin{array}[c]{c}e_{1}+l_{1}+1\\ e_{a}+l_{a}+1;e_{b}+l_{b}+1\end{array}\right)\right]^{1/2}.$ (19)
An additional component in the eq.(14) is the 9j symbol, denoted by the curly brackets. The 9j symbol in the expression has eight summation indices, although only three are independent. The described expression has proven to be useful in the ab initio nuclear structure calculations [17, 18, 19, 20, 21]. It is well-balanced and gives fast and precise results. However, the calculation of the 9j coefficient is still expensive, so an alternative expression that could producen nice results would be useful.
3 An alternative expression: HOB in $SU(3)$ basis
In this section, we formulate the motivation for the alternative expression in the $SU(3)$ basis. In quantum mechanics, the wavefunction is a scalar product of two vectors in different bases in the Hilbert space. For two HOs we can write the wavefunction as
$\varphi_{e_{1}l_{1},e_{2}l_{2}:LM}(\boldsymbol{r}_{1},\boldsymbol{r}_{2})\equiv\{\phi_{e_{1}l_{1}}(\boldsymbol{r}_{1})\otimes\phi_{e_{2}l_{2}}(\boldsymbol{r}_{2})\}_{LM}\equiv\left\langle\boldsymbol{r}_{1},\boldsymbol{r}_{2}|e_{1}l_{1},e_{2}l_{2}:LM\right\rangle.$ (20)
The intrinsic symmetry of a quantum three-dimensional HO is $U(3).$ If we have two three-dimensional HOs then the group describing the system is $U(6).$ Lets mark the symmetric irreducible representation (irrep) as $[E]$. $E$ is the total HO energy for the system, ergo $E=e_{1}+e_{2}$. The basis vector as
$|e_{1}l_{1}e_{2}l_{2}:LM\rangle.$ (21)
This is the conventional basis for this problem we introduced in the previous section. The group chain for this basis is
$\begin{array}[c]{ccccc}U(6)&\supset&U(3)&+&U(3)\\ &&\cup&&\cup\\ &&SO(3)&&SO(3)\end{array}.$ (22)
Here the quantum numbers for the irrep labelling are $e_{1},$ $e_{2}$ and $l_{1},l_{2}.$ The angular momenta $l_{1}$ and $l_{2}$ are coupled to the total angular momentum $L.$ Quantum numbers $e_{1},$ $e_{2}$ are the energy of the HOs. However, we can try to use a different group chain for the $U(6)$, namely
$\begin{array}[c]{ccccc}U(6)&\supset&U(3)&\times&U(2)\\ &&\cup&&\cup\\ &&SO(3)&&SO(2)\end{array}.$ (23)
For this group chain the basis is
$|e_{1}e_{2}[E_{1}E_{2}]\alpha LM\rangle.$ (24)
The quantum numbers $e_{1},$ $e_{2}$ are the energy of the HOs, $[E_{1}E_{2}]$ label the irreps of $U_{3}$, $L$ is the angular momentum and $M$ is the projection of the angular momentum along the z-axis. $\alpha$ is the index of recurrence of $L$ in the irrep $[E_{1}E_{2}]$
The basis in eq.(24) can also be written as
$|e_{1}e_{2}J\alpha LM\rangle$ (25)
if we denote $\frac{E_{1}-E_{2}}{2}$ as $J$. The group chain $SU(3)\supset SO(3)\supset SO(2)$ is a non-canonical subgroup chain.
A very important tool for coupling two $SU(3)$ representations are the Clebsch-Gordan coefficients, or their isofactors. These coefficients emerge when we do a reduction of a product of representations of $SU(3)$ into the irreps of $SO(3).$ The correct combination of the quantum states from both groups must be ensured. For the two irreducible representations $(\lambda_{1},\mu_{1})$ and $(\lambda_{2},\mu_{2})$ of $SU(3)$ the decomposition is
$\{\lambda_{1},\mu_{1}\}\otimes\{\lambda_{2},\mu_{2}\}=\sum_{\lambda^{\prime},\mu^{\prime}}C_{(\lambda^{\prime},\mu^{\prime})}^{(\lambda_{1},\mu_{1}),(\lambda_{2},\mu_{2})}.$ (26)
The irreps of $SO(3)$ are marked as $\{\lambda^{\prime},\mu^{\prime}\}$    and the Clebsch-Gordan coefficients $C_{(\lambda^{\prime},\mu^{\prime})}^{(\lambda_{1},\mu_{1}),(\lambda_{2},\mu_{2})}$ are the isofactors. In agreement with Racah lemma [22] the matrix elements
$\langle(e_{1}l_{1},e_{2}l_{2})L|(e_{1},e_{2})[E_{1}E_{2}]\alpha L\rangle\equiv\langle e_{1}l_{1},e_{2}l_{2}:L|e_{1}e_{2}J\alpha L\rangle$ (27)
that transform the vectors from one basis to the vectors of the second basis are isofactors of Clebsch-Gordan coefficients of the group $SU(3).$ These coefficients are used for the non-canonical embedding of the irreps. They can also be denoted as $\left[\begin{array}[c]{ccc}e_{1}&e_{2}&j\alpha\\ l_{1}&l_{2}&L\end{array}\right]$ or $C_{l_{1}l_{2}L}^{e_{1}e_{2},j\alpha}$ or $\langle l_{1}l_{2}|\alpha j\rangle^{e_{1}e_{2}L}$. We will use the last notation.
4 The Talmi-Moshinsky transformation in $SU(3)$ representation
In the first section, we presented the Talmi-Moshinsky transformation as the two-dimensional matrix
$\langle\boldsymbol{r}_{1}^{\prime},\boldsymbol{r}_{2}^{\prime}|\ \boldsymbol{r}_{1},\boldsymbol{r}_{2}\rangle_{d}=\left(\begin{tabular}[]{ll}$\sqrt{\frac{d}{1+d}}$&$\phantom{-}\sqrt{\frac{1}{1+d}}$\\ $\sqrt{\frac{1}{1+d}}$&$-\sqrt{\frac{d}{1+d}}$\end{tabular}\right).$ (28)
This matrix has a determinant $-1$ and is a representation of the group $O(2)$. Since $O(2)\supset SO(2)$, we rewrite the Talmi-Moshinsky (TM) transformation as a representation of the $R$$\times$ $SO(2)$, where $R$ is reflection part of the transformation
d+1d 1+1d
1+1d -d+1d
=(1001)×(d1+d11+d11+dd1+d)
d+1d 1+1d
1+1d -d+1d
[…]
Further, we will investigate the TM transformation without the reflection. We write the TM as a general linear transformation matrix
$\widehat{a}=\left(\begin{tabular}[]{ll}$a_{1}$&$a_{3}$\\ $a_{2}$&$a_{4}$\\ \end{tabular}\right)$ (30)
and impose the properties of unitarity and unimodularity- $\widehat{a}\widehat{a}^{+}=1$ and $\det\widehat{a}=1.$ Immediately, we see that $a_{2}=-a_{3}^{\ast}$ and $a_{4}=a_{1}^{\ast}$ producing
$\widehat{a}=\left(\begin{tabular}[]{ll}$\phantom{-}a_{1}$&$a_{3}$\\ $-a_{3}^{\ast}$&$a_{1}^{\ast}$\\ \end{tabular}\right),$ (31)
parametrized by the Cayley-Klein parameters [23]. The matrix (31) is the representation of the $SU(2),$ which is the double covering group of $SO(2).$ We can specify the matrix (31) elements as
cos(θ) sin(θ)
-sin(θ) cos(θ)
cos(θ) sin(θ)
-sin(θ) cos(θ)
[…]
In the matrix (32) TM transformation is parametrized with an angle $\theta,$ which is connected with $d$ in a straightforward way
$\frac{1}{1+d}=\sin^{2}\theta,$ (33)
$\frac{1}{\sin^{2}\theta}=1+d,$ (34)
$d=\frac{1}{\sin^{2}\theta}-1=\frac{\sin^{2}\theta+\cos^{2}\theta-\sin^{2}\theta}{\sin^{2}\theta}=\cot^{2}\theta.$ (35)
We need to find a representation of the transformation matrix (32) in a $SU(3)$ basis. This orthogonal transformation must transform the function of the coupled $SU(3)$ irreps $e_{1}$ and $e_{2}$, depending on the variables $\boldsymbol{r}_{1}$ and $\boldsymbol{r}_{2}$. The action of the transformation would result in a function depending on the transformed variables $\boldsymbol{r}_{1}^{\prime}$ and $\boldsymbol{r}_{2}^{\prime}.$ We write the wavefunction of two coupled HOs in $SU(3)$ as $\Psi(e_{1}e_{2}[E_{1}E_{2}]\gamma|\boldsymbol{r}_{1},\boldsymbol{r}_{2})$ or in a bra-ket notation as $\langle\boldsymbol{r}_{1},\boldsymbol{r}_{2}|e_{1}e_{2}[E_{1}E_{2}]\gamma\rangle,$ where $\gamma=\alpha LM.$
If the ket vector is transformed by some operator $T$
$\langle\phi|T|\psi\rangle,$ (36)
then to keep the scalar product invariant, we need to act on a bra vector with a some sort of operator $X$ in a such way that
$X\langle\phi|T|\psi\rangle=\langle\phi|\psi\rangle.$ (37)
This procedure implies that the operator $X$ must satisfy the equality
$X\langle\phi|=\langle\phi|T^{\dagger}.$ (38)
Thus, the operator $X$ is the Hermitian conjugate of the operator $T$ and can be denoted as $T^{\dagger}$. We follow this outline and insert the transformation $\hat{a}$ and its complex conjugate $\hat{a}^{\dagger}=\hat{a}^{-1}$ in our bra-ket $\langle\boldsymbol{r}_{1},\boldsymbol{r}_{2}\left|\hat{a}^{-1}\hat{a}\right|e_{1}e_{2}[E_{1}E_{2}]\gamma\rangle$. We introduce the intermediate summation of $\frac{(e_{1}^{\prime}-e_{2}^{\prime})}{2}$ and integrate out the operator $a^{-1}$ using the completeness relation for discrete and continuous spectra
[…]
×$\times{\textstyle\sum\limits_{(e_{1}^{\prime}-e_{2}^{\prime})/2}}\langle\boldsymbol{r}_{1}^{\prime},\boldsymbol{r}_{2}^{\prime}|e_{1}^{\prime}e_{2}^{\prime}[E_{1}E_{2}]\gamma\rangle\langle e_{1}^{\prime}e_{2}^{\prime}[E_{1}E_{2}]\gamma|\hat{a}|e_{1}e_{2}[E_{1}E_{2}]\gamma\rangle$
$={\textstyle\sum\limits_{(e_{1}^{\prime}-e_{2}^{\prime})/2}}\langle\boldsymbol{r}_{1}^{\prime},\boldsymbol{r}_{2}^{\prime}|e_{1}^{\prime}e_{2}^{\prime}[E_{1}E_{2}]\gamma\rangle\langle e_{1}^{\prime}e_{2}^{\prime}[E_{1}E_{2}]\gamma|\hat{a}|e_{1}e_{2}[E_{1}E_{2}]\gamma\rangle$
$=\Delta_{0}^{E_{2}}{\textstyle\sum\limits_{(e_{1}^{\prime}-e_{2}^{\prime})/2}}\Psi(e_{1}^{\prime}e_{2}^{\prime}[E_{1}E_{2}]\gamma|\boldsymbol{r}_{1}^{\prime},\boldsymbol{r}_{2}^{\prime})d_{(e_{1}^{\prime}-e_{2}^{\prime})/2,(e_{1}-e_{2})/2}^{(E_{1}-E_{2})/2}(\theta).$ (39)
The action of the operator $\hat{a}$ results in a Wigner d-matrix of the irreducible representation of the group $SU(2)$, which is a subgroup of $SU(3)$. It is an orthogonal transformation that transforms a function of the coupled $SU(3)$ irreps $e_{1}$ and $e_{2}$ depending on the variables $\boldsymbol{r}_{1}$ and $\boldsymbol{r}_{2}$ into a function depending on the variables $\boldsymbol{r}_{1}^{\prime}$ and $\boldsymbol{r}_{2}^{\prime}$. The Wigner d-matrix depends on the same angle $\theta$. We change the notation a little bit, folowing the eq.(25), namely
$J=\frac{(E_{1}-E_{2})}{2};M=\frac{(e_{1}-e_{2})}{2};M^{\prime}=\frac{(e_{1}^{\prime}-e_{2}^{\prime})}{2}.$ (40)
This results in more common expression for eq.(4)
$\langle\boldsymbol{r}_{1},\boldsymbol{r}_{2}\left|\hat{a}^{-1}\hat{a}\right|e_{1}e_{2}[E_{1}E_{2}]\gamma\rangle=\Delta_{0}^{E_{2}}{\textstyle\sum\limits_{m^{\prime}}}\Psi(e_{1}^{\prime}e_{2}^{\prime}[E_{1}E_{2}]\gamma|\boldsymbol{r}_{1}^{\prime},\boldsymbol{r}_{2}^{\prime})d_{M,M^{\prime}}^{J}(\theta).$ (41)
The different expressions for the Wigner d-matrix are well known. For example, we can use the expression from [24].
$d_{M,M^{\prime}}^{J}(\theta)=\sum_{k}\frac{\sqrt{(J+M)!(J-M)!(J+M^{\prime})!(J-M^{\prime})!}}{k!(J-M-k)!(J+M^{\prime}-k)!(k+M-M^{\prime})!}$
×$\times a_{1}^{J+M^{\prime}-k}a_{2}^{k}a_{3}^{k+M-M^{\prime}}a_{4}^{J-M-k}.$ (42)
This is a general expression, which is parametrized by a matrix (30). From bounded conditions of the matrix (31) we get the connections
[…]
$a_{4}$ $=-\cos(\theta),$ (44)
$a_{2}$ $=a_{3}=\sin(\theta).$ (45)
Inserting them into the eq.(42) results in an expression
$d_{M,M^{\prime}}^{J}(\theta)$ $=\sum_{k}\frac{\sqrt{(J+M)!(J-M)!(J+M^{\prime})!(J-M^{\prime})!}}{k!(J-M-k)!(J+M^{\prime}-k)!(k+M-M^{\prime})!}$
×$\times-\cos(\theta)^{J+M^{\prime}-k}\sin(\theta)^{k}\sin(\theta)^{k+M-M^{\prime}}\cos(\theta)^{J-M-k}.$ (46)
We rearange this expression to make it computationally more feasible by using $\Gamma$ function instead of factorials and pulling out terms of equation independent from the summation
$d_{M,M^{\prime}}^{J}(\theta)=\sqrt{\Gamma(J+M+1)\Gamma(J-M+1)\Gamma(J+M^{\prime}+1)\Gamma(J-M^{\prime}+1)}$
×$\times\sum_{k}\frac{-\cos(\theta)^{2J+M^{\prime}-M-2k}\sin(\theta)^{2k+M-M^{\prime}}}{\Gamma(k+1)\Gamma(J-M-k+1)\Gamma(J+M^{\prime}-k+1)\Gamma(k+M-M^{\prime}+1)}.$ (47)
The momentum $J$ belongs in the interval $|m|\leqslant J\leqslant J_{\max},$ where the $J_{\max}$ is equal to $\frac{E}{2}.$ Each $J$ has its own values of $L$ recurrence index $\alpha$. The recurrence index $\alpha$, also known as the degeneracy index, in a $SU(3)$ irrep can be determined by the Racah $d_{(\lambda,\mu)}^{L}$ formula [25]. The index $(\lambda,\mu)$ is yet another marking of the $SU(3)$ irrep, which is connected to the notation presented before. Since we use quantum numbers $E\alpha J$, we rewrite the Racah formula
$d_{(\lambda,\mu)}^{L}=\left[\frac{\lambda+\mu-L+2}{2}\right]_{\geq 0}-\left[\frac{\lambda-L+1}{2}\right]_{\geq 0}-\left[\frac{\mu-L+1}{2}\right]_{\geq 0},$ (48)
where $[\dots]_{\geq 0}$ is zero for any negative number and gives the positive integer part otherwise. This results in the equation
$d_{(\lambda,\mu)}^{L}=\alpha_{EJ}^{L}=\left[\frac{E_{1}-L+2}{2}\right]_{\geq 0}-\left[\frac{2J-L+1}{2}\right]_{\geq 0}-\left[\frac{E_{2}-L+1}{2}\right]_{\geq 0}$
$=\left[\frac{E/2+J-L+2}{2}\right]_{\geq 0}-\left[\frac{2J-L+1}{2}\right]_{\geq 0}-\left[\frac{E/2-J-L+1}{2}\right]_{\geq 0}.$ (49)
As a result we see that for the given $EJL$ values the index $\alpha$ runs in the interval $1\leqslant\alpha\leqslant\alpha_{EJ}^{L}=\alpha_{0}$. This is a crucial information for the required $SU(3)$ Clebsch-Gordan coefficient isofactor calculation. If $\alpha_{EJ}^{L}$ equals to zero, then the $SU(3)$ irrep $EJL$ does not exist. We can calculate the dimension of the irrep using the expression
$\dim(\lambda,\mu)=(1+\lambda)(1+\mu)\left(1+\frac{\lambda+\mu}{2}\right)$ (50)
from [26], which we also rewrite in the $E\alpha J$ notation as
$\dim(Ej)=(1+2J)(1+E/2-J)\left(1+\frac{E/2+J}{2}\right).$ (51)
One can check that we indeed have
$d_{(\lambda,\mu)}=\sum_{L}(2L+1)d_{(\lambda,\mu)}^{L}.$ (52)
We calculate the reflection $R$ from the matrix
$a=\frac{1}{\Delta_{0}}\left(\begin{tabular}[]{ll}$a_{1}$&$a_{3}$\\ $a_{2}$&$a_{4}$\\ \end{tabular}\right)$ (53)
and the representation $d_{M,M^{\prime}}^{J}(\theta).$ When the matrix is given a form ($1$&$\phantom{-}0$\\ $0$&$-1$\\ \end{tabular}\right),$ we see that its representation is $d_{M,M^{\prime}}^{J}=(-1)^{J+M}\delta(M,M^{\prime}).$
5 The calculation of the HOB
In this section, we gather the presented knowledge into a new approach for the calculation of the HOB. HOB is defined as $\langle e_{1}l_{1},e_{2}l_{2}:L|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:L\rangle_{d}$, or $\langle e_{1}l_{1},e_{2}l_{2}:L|T(d)|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:L\rangle.$ We transform the $SO(3)$ basis into the $SU(3)$ basis by using the Clebsh-Gordan isofactors and summing though the intermediate states
$\langle e_{1}l_{1},e_{2}l_{2}:L|T(d)|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:L\rangle={\textstyle\sum\limits_{E_{1}E_{2}\alpha E_{1}^{\prime}E_{2}^{\prime}\alpha^{\prime}}}\langle e_{1}l_{1},e_{2}l_{2}:L|e_{1},e_{2}[E_{1}E_{2}]\alpha:L\rangle$
×$\times\langle e_{1},e_{2}[E_{1}E_{2}]\alpha:L|T(d)|e_{1}^{\prime},e_{2}^{\prime}[E_{1}^{\prime}E_{2}^{\prime}]\alpha^{\prime}:L\rangle\langle e_{1}^{\prime},e_{2}^{\prime}[E_{1}^{\prime}E_{2}^{\prime}]\alpha^{\prime}|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:L\rangle$
${\textstyle\sum\limits_{JM\alpha J^{\prime}M^{\prime}\alpha^{\prime}}}\langle e_{1}l_{1},e_{2}l_{2}:L|e_{1},e_{2}J\alpha:L\rangle\langle e_{1},e_{2}JM\alpha:L|T(d)|e_{1}^{\prime},e_{2}^{\prime}J^{\prime}M^{\prime}\alpha^{\prime}:L\rangle$
×$\times\langle e_{1}^{\prime},e_{2}^{\prime}J^{\prime}\alpha^{\prime}:L|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:L\rangle\delta_{J,J^{\prime}}\text{ }\delta_{\alpha,\alpha^{\prime}}.$ (54)
We make this expression more compact rewriting in more favorable notation.
$\langle e_{1}l_{1},e_{2}l_{2}:L|T(d)|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:L\rangle$
$={\textstyle\sum\limits_{JM\alpha J^{\prime}M^{\prime}\alpha^{\prime}}}\langle l_{1}l_{2}|J\alpha\rangle^{e_{1}e_{2}L}d_{M,M^{\prime}}^{\ J}(d)\langle J^{\prime}\alpha^{\prime}|l_{1}^{\prime}l_{2}^{\prime}\rangle^{e_{1}^{\prime}e_{2}^{\prime}L}\delta_{\alpha,\alpha^{\prime}}\delta_{J,J^{\prime}}.$ (55)
By taking into account the $\delta$ functions we get the expression
$\langle e_{1}l_{1}e_{2}l_{2}L|T(d)|e_{1}^{\prime}l_{1}^{\prime}e_{2}^{\prime}l_{2}^{\prime}L\rangle={\textstyle\sum\limits_{JMaM^{\prime}}}\langle l_{1}l_{2}|J\alpha\rangle^{e_{1}e_{2}L}d_{M,M^{\prime}}^{\ J}(d)\langle J\alpha|l_{1}^{\prime}l_{2}^{\prime}\rangle^{e_{1}^{\prime}e_{2}^{\prime}L}.$ (56)
We further simplify this expression by using the relations
$e_{1}-e_{2}=2M;$ (57)
$e_{1}^{\prime}-e_{2}^{\prime}=2M^{\prime};$ (58)
$e_{1}+e_{2}=e_{1}^{\prime}+e_{2}^{\prime}=E.$ (59)
This results in the equation
$\langle e_{1}l_{1}e_{2}l_{2}L|T(d)|e_{1}^{\prime}l_{1}^{\prime}e_{2}^{\prime}l_{2}^{\prime}L\rangle=$
${\textstyle\sum\limits_{J=\max(|M|,|M^{\prime}|)}^{E/2}}d_{M,M^{\prime}}^{\ J}(d){\textstyle\sum\limits_{\alpha=1}^{\alpha_{0}}}\langle l_{1}l_{2}|J\alpha\rangle^{e_{1}e_{2}L}\langle l_{1}^{\prime}l_{2}^{\prime}|J\alpha\rangle^{e_{1}^{\prime}e_{2}^{\prime}L}.$ (60)
We are left with two summations of permitted $J$ and $\alpha$ values with bounds $|M|\ \leq J\leq E/2$ and $|M^{\prime}|\ \leq J\leq E/2$. The multiplicity label $\alpha$ enumerates the reccuring $SO(3)$ irrep $L$ in the $SU(3)$ irrep $J$. The CG coefficients in the SO(3) bases are constructed subsequently from the overlaps between the SU(2) and SO(3) bases.
6 Computation of $SU(3)$ HOB
The computation of the eq.(60) is straightforward. For testing we used the matrix multiplication version, matrix element code, and compared the results to a HOB calculation code from [5]. We employed Fortran90 language. For the $SU(3)\supset SO(3)$ CG coefficient calculation we use the computer code created by C. Bahri, D.J. Rowe, and J.P. Draayer [27]. The eq.(4) is computed using the standard Fortran Gamma function. The calculation effort of a Wigner D matrix element is negligible and all the heavy lifting is done by the CG coefficients. We used a GNU Fortran compiler and the profiling was done using gprof. We checked the calculation time for HO E=[0,20] with all the L’s=[0,E] and found the CG coefficient computation takes 82.46 % of the time in function for matrix element computation. The d-matrix element calculation takes only 0.51 % of all the calculation. The rest are auxiliary functions. We calculated the HOB matrix for different E and L which results in dense matrices. The matrix for a given E would have a block-diagonal structure depending on L. This structure allows to build matrix for a given HO E by calculating only blocks. We sum the time and the error for all the values of L for a given HO E.
E $\delta{(MatMult)}$ $\delta{(HOB)}$ $\delta{(SU3HOB)}$
0 0.100E-04 0.100E-05 0.800E-05
1 0.500E-05 0.100E-05 0.160E-04
2 0.120E-04 0.400E-05 0.460E-04
3 0.280E-04 0.150E-04 0.172E-03
4 0.760E-04 0.580E-04 0.643E-03
5 0.155E-03 0.151E-03 0.236E-02
6 0.421E-03 0.491E-03 0.737E-02
7 0.811E-03 0.117E-02 0.216E-01
8 0.151E-02 0.264E-02 0.509E-01
9 0.272E-02 0.533E-02 0.115E+00
10 0.490E-02 0.113E-01 0.265E+00
11 0.851E-02 0.247E-01 0.569E+00
12 0.152E-01 0.457E-01 0.118E+01
13 0.239E-01 0.839E-01 0.237E+01
14 0.393E-01 0.151E+00 0.460E+01
15 0.651E-01 0.291E+00 0.874E+01
16 0.102E+00 0.469E+00 0.163E+02
17 0.153E+00 0.799E+00 0.288E+02
18 0.226E+00 0.131E+01 0.494E+02
19 0.325E+00 0.209E+01 0.824E+02
20 0.491E+00 0.346E+01 0.137E+03
21 0.701E+00 0.548E+01 0.222E+03
22 0.101E+01 0.851E+01 0.353E+03
23 0.141E+01 0.132E+02 0.555E+03
Table 1: Calculation time in seconds of HOB block matrices for a given HO Energy quanta. E is the HO Energy, Matmult- Matrix multiplication version of the code, HOB code [5], SU3HOB- matrix element code.
In the table (1) we see that the calculation of the MatMult method is faster by a factor of ten, than the HOB code from [5]. The SU3HOB matrix element code allows the computation of the HOB matrix by a factor of ten slower.
For the error calculation, we use the fact that the calculated matrices must be orthonormal, that is $HOBxHOB^{T}=I.$ Thus we can calculate the deviation from the unitary matrix. We norm the result using the dimension of the matrix to see the deterioration of the matrix as for larger matrices we would have more error. We also sum for a given HO E. The normed error values for MatMult and SU3HOB are the same. The results are presented in table (2).
E $\delta{(SU3HOB)}$ $\delta{(HOB)}$
0 0.143E-12 0.000E+00
1 0.141E-12 0.444E-15
2 0.141E-12 0.460E-15
3 0.142E-12 0.791E-15
4 0.143E-12 0.178E-14
5 0.143E-12 0.372E-14
6 0.145E-12 0.668E-14
7 0.146E-12 0.166E-13
8 0.162E-12 0.649E-13
9 0.185E-12 0.153E-12
10 0.424E-12 0.746E-12
11 0.878E-12 0.231E-11
12 0.201E-11 0.632E-11
13 0.537E-11 0.327E-10
14 0.269E-10 0.106E-09
15 0.710E-10 0.505E-09
16 0.262E-09 0.134E-08
17 0.713E-09 0.503E-08
18 0.189E-08 0.190E-07
19 0.790E-08 0.683E-07
20 0.328E-07 0.273E-06
21 0.155E-06 0.924E-06
22 0.687E-06 0.331E-05
23 0.279E-05 0.133E-04
Table 2: Error of the HOB matrices calculated by HOB and SU3HOB for a given HO Energy quanta. E is the HO Energy, HOB code from [5], SU3HOB- matrix element code.
We see that for small HO E up until E=12, the HOB code is more accurate, however for larger E, the SU3HOB takes the lead. More detailed analysis of data shows that for larger matrices where L values are up until approximately $\frac{E}{2},$ the HOB code deteriotes faster. However, for larger Ls, the SU3HOB code gives worse results (see table (3)). This can be attributed to the SU3 Clebsh-Gordan code, which has trouble coupling for the larger Ls. We checked the errors for the Clebsh-Gordan matrix and the d-matrix used for matrix multiplication code following the same condition of orthonormality. The results are presented in table (4). We see a clear deterioration of the CG matrix, while the d matrix is pretty stable.
E L Dimension $\delta{(HOB)}$ $\delta{(SU3HOB)}$
23 1 156 0.176E-04 0.118E-08
23 2 132 0.124E-04 0.909E-09
23 3 264 0.118E-03 0.337E-08
23 4 220 0.110E-03 0.304E-08
23 5 330 0.540E-03 0.442E-07
23 6 270 0.480E-03 0.398E-07
23 7 360 0.250E-02 0.155E-06
23 8 288 0.159E-02 0.138E-06
23 9 360 0.125E-01 0.533E-06
23 10 280 0.472E-02 0.475E-06
23 11 336 0.233E-01 0.118E-05
23 12 252 0.414E-02 0.105E-05
23 13 294 0.101E-01 0.234E-05
23 14 210 0.881E-03 0.204E-05
23 15 240 0.190E-02 0.155E-04
23 16 160 0.955E-04 0.136E-04
23 17 180 0.327E-04 0.878E-04
23 18 108 0.214E-05 0.744E-04
23 19 120 0.428E-06 0.431E-03
23 20 60 0.166E-07 0.372E-03
23 21 66 0.101E-08 0.297E-02
23 22 22 0.450E-10 0.256E-02
23 23 24 0.842E-11 0.668E-02
Table 3: Error of the HOB matrices calculated by 3HOB and SU3HOB for a given HO Energy quanta. E is the HO Energy, 3HOB code from [5], SU3HOB- matrix element code of derived expression.
E $\delta{(CGMat)}$ $\delta{(Djmm^{\prime})}$
1 0.141E-12 0.444E-15
2 0.493E-12 0.266E-14
3 0.850E-12 0.344E-14
4 0.186E-11 0.133E-13
5 0.280E-11 0.231E-13
6 0.496E-11 0.575E-13
7 0.701E-11 0.754E-13
8 0.114E-10 0.149E-12
9 0.165E-10 0.200E-12
10 0.521E-10 0.372E-12
11 0.141E-09 0.483E-12
12 0.411E-09 0.901E-12
13 0.153E-08 0.126E-11
14 0.935E-08 0.242E-11
15 0.293E-07 0.373E-11
16 0.133E-06 0.652E-11
17 0.422E-06 0.952E-11
18 0.136E-05 0.152E-10
19 0.704E-05 0.219E-10
20 0.371E-04 0.342E-10
21 0.208E-03 0.495E-10
22 0.111E-02 0.746E-10
23 0.523E-02 0.111E-09
Table 4: Error of the Clebsh-Gordan matrix compared to Wigner d-matrix for a given HO Energy quanta.
7 Conclusions
We have introduced a new approach for the harmonic oscillator brackets using the SU(3) scheme. The SU(3) scheme allows a simple and elegant expression for the calculation of the Talmi-Moshinsky transformation representation in the HO basis. We compared our calculation time and errors to widely used code [5]. While our matrix element code lags behind the HOB code, we get more accuracy. When the whole HOB matrix is needed for given HO E and L, the matrix multiplication approach is superior from both accuracy and time consumption points of view. All the heavy lifting in the calculation of the SU3HOB is done by the Clebsch-Gordan coefficient calculation code [27]. Optimization of the said coefficient calculation would allow us to get faster and maybe more accurate results.
Another important aspect is that in this scheme we can calculate the Wigner d-matrix with different kinds of d parameters for various Jacobi coordinate transformations and calculate the transformation between the SO(3) and SU(3) basis only once. This approach allows an exciting opportunity to formulate the ab initio no-core shell model in the unitary scheme using Jacobi coordinates.
Acknowledgements
This project has received funding from the Research Council of Lithuania (LMTLT), agreement No S-PD-22-9
The authors gratefully acknowledge the computing time granted on the supercomputer JURECA [28] at Forschungszentrum Juelich.

### v2

@@
 in [5].
-Here $d$ is a nonnegative real number.
+Here $d$ is a nonnegative real number, parameterizing the Talmi-Moshinsky transformation (see eq.(5)).
 The variables in the brackets are arranged in a fixed way and the orthogonal transformation of the coordinates $\boldsymbol{r}_{1},\boldsymbol{r}_{2}$ is written as the two-dimensional matrix
 $\langle\boldsymbol{r}_{1}^{\prime},\boldsymbol{r}_{2}^{\prime}|\ \boldsymbol{r}_{1},\boldsymbol{r}_{2}\rangle_{d}=\left(\begin{tabular}[]{ll}$\sqrt{\frac{d}{1+d}}$&$\phantom{-}\sqrt{\frac{1}{1+d}}$\\ $\sqrt{\frac{1}{1+d}}$&$-\sqrt{\frac{d}{1+d}}$\end{tabular}\right).$ (5)
-This two-dimensional matrix is called the Talmi-Moshinsky transformation. It is parametrized by the variable $d.$ In [5] the original expression for HOBs derived by B.Buck and A.C.Merchant [16]
+This two-dimensional matrix is called the Talmi-Moshinsky transformation. In [5] the original expression for HOBs derived by B.Buck and A.C.Merchant [16]
 $\left\langle EL,el:\Lambda\right.\left|e_{1}l_{1},e_{2}l_{2}:\Lambda\right\rangle_{d}$
@@
 It is well-balanced and gives fast and precise results.
-However, the calculation of the 9j coefficient is still expensive, so an alternative expression that could producen nice results would be useful.
+However, the calculation of the 9j coefficient is still expensive, so an alternative expression that could produce nice results would be useful.
 3 An alternative expression: HOB in $SU(3)$ basis
@@
 Since $O(2)\supset SO(2)$, we rewrite the Talmi-Moshinsky (TM) transformation as a representation of the $R$$\times$ $SO(2)$, where $R$ is reflection part of the transformation
-d+1d 1+1d
-1+1d -d+1d
-=(1001)×(d1+d11+d11+dd1+d)
-d+1d 1+1d
-1+1d -d+1d
-[…]
-[…]
+($\sqrt{\frac{d}{1+d}}$&$\phantom{-}\sqrt{\frac{1}{1+d}}$\\ $\sqrt{\frac{1}{1+d}}$&$-\sqrt{\frac{d}{1+d}}$\\ \end{tabular}\right)=\left(\begin{tabular}[]{ll}$1$&$\phantom{-}0$\\ $0$&$-1$\\ \end{tabular}\right)\times\left(\begin{tabular}[]{ll}$\phantom{-}\sqrt{\frac{d}{1+d}}$&$\sqrt{\frac{1}{1+d}}$\\ $-\sqrt{\frac{1}{1+d}}$&$\sqrt{\frac{d}{1+d}}$\\ \end{tabular}\right).$ (29)
 Further, we will investigate the TM transformation without the reflection. We write the TM as a general linear transformation matrix
@@
 parametrized by the Cayley-Klein parameters [23]. The matrix (31) is the representation of the $SU(2),$ which is the double covering group of $SO(2).$ We can specify the matrix (31) elements as
-cos(θ) sin(θ)
--sin(θ) cos(θ)
-cos(θ) sin(θ)
--sin(θ) cos(θ)
-[…]
+($\phantom{-}\cos(\theta)$&$\sin(\theta)$\\ $-\sin(\theta)$&$\cos(\theta)$\\ \end{tabular}\right).$ (32)
 In the matrix (32) TM transformation is parametrized with an angle $\theta,$ which is connected with $d$ in a straightforward way
@@
 The Wigner d-matrix depends on the same angle $\theta$.
-We change the notation a little bit, folowing the eq.(25), namely
+We change the notation a little bit, following the eq.(25), namely
 $J=\frac{(E_{1}-E_{2})}{2};M=\frac{(e_{1}-e_{2})}{2};M^{\prime}=\frac{(e_{1}^{\prime}-e_{2}^{\prime})}{2}.$ (40)
@@
 ×$\times-\cos(\theta)^{J+M^{\prime}-k}\sin(\theta)^{k}\sin(\theta)^{k+M-M^{\prime}}\cos(\theta)^{J-M-k}.$ (46)
-We rearange this expression to make it computationally more feasible by using $\Gamma$ function instead of factorials and pulling out terms of equation independent from the summation
+We rearrange this expression to make it computationally more feasible by using $\Gamma$ function instead of factorials and pulling out terms of equation independent from the summation
 $d_{M,M^{\prime}}^{J}(\theta)=\sqrt{\Gamma(J+M+1)\Gamma(J-M+1)\Gamma(J+M^{\prime}+1)\Gamma(J-M^{\prime}+1)}$
@@
 $\dim(\lambda,\mu)=(1+\lambda)(1+\mu)\left(1+\frac{\lambda+\mu}{2}\right)$ (50)
-from [26], which we also rewrite in the $E\alpha J$ notation as
+from [26]. which we also rewrite in the $E\alpha J$ notation as
 $\dim(Ej)=(1+2J)(1+E/2-J)\left(1+\frac{E/2+J}{2}\right).$ (51)
@@
 In this section, we gather the presented knowledge into a new approach for the calculation of the HOB.
-HOB is defined as $\langle e_{1}l_{1},e_{2}l_{2}:L|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:L\rangle_{d}$, or $\langle e_{1}l_{1},e_{2}l_{2}:L|T(d)|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:L\rangle.$ We transform the $SO(3)$ basis into the $SU(3)$ basis by using the Clebsh-Gordan isofactors and summing though the intermediate states
+HOB is defined as $\langle e_{1}l_{1},e_{2}l_{2}:L|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:L\rangle_{d}$, or $\langle e_{1}l_{1},e_{2}l_{2}:L|T(d)|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:L\rangle.$ We transform the $SO(3)$ basis into the $SU(3)$ basis by using the Clebsch-Gordan isofactors and summing though the intermediate states
 $\langle e_{1}l_{1},e_{2}l_{2}:L|T(d)|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:L\rangle={\textstyle\sum\limits_{E_{1}E_{2}\alpha E_{1}^{\prime}E_{2}^{\prime}\alpha^{\prime}}}\langle e_{1}l_{1},e_{2}l_{2}:L|e_{1},e_{2}[E_{1}E_{2}]\alpha:L\rangle$
@@
 The multiplicity label $\alpha$ enumerates the reccuring $SO(3)$ irrep $L$ in the $SU(3)$ irrep $J$.
+The label $\alpha$ is the same as the label $K$ introduced by J.
+P.
+Elliott in [27].
 The CG coefficients in the SO(3) bases are constructed subsequently from the overlaps between the SU(2) and SO(3) bases.
@@
 However, for larger Ls, the SU3HOB code gives worse results (see table (3)).
-This can be attributed to the SU3 Clebsh-Gordan code, which has trouble coupling for the larger Ls.
-We checked the errors for the Clebsh-Gordan matrix and the d-matrix used for matrix multiplication code following the same condition of orthonormality.
+This can be attributed to the SU3 Clebsch-Gordan code, which has trouble coupling for the larger Ls.
+We checked the errors for the Clebsch-Gordan matrix and the d-matrix used for matrix multiplication code following the same condition of orthonormality.
 The results are presented in table (4).
@@
 23 0.523E-02 0.111E-09
-Table 4: Error of the Clebsh-Gordan matrix compared to Wigner d-matrix for a given HO Energy quanta.
+Table 4: Error of the Clebsch-Gordan matrix compared to Wigner d-matrix for a given HO Energy quanta.
 7 Conclusions